For the construction of a customs warehouse, an investment of 10,000 is required for two consecutive years. Then, the for next 3 years the warehouse will bring 6500 in profit per year. Calculate NPV and IRR and evaluate the project. The investor has an alternative to invest money at 15% per annum.
1
Expert's answer
2020-12-29T17:59:01-0500



If the initial investment each year is 10,000 in the first two years, then the project does not pay off, the NPV and IRR are less than zero. By reducing the initial investment to 5,000 per year, the project pays off


To calculate the NPV, you must:

Make a forecast schedule of cash flows for the investment project in the context of periods. Cash flows should include both income (inflows of funds) and expenses (investments made and other project implementation costs).

Determine the size of the discount rate. In fact, the discount rate reflects the marginal rate of the investor's cost of capital. For example, if the bank's borrowed funds are used for investment, the discount rate will be the effective interest rate on the loan. If the investor's own funds are used, then the discount rate can be taken as the interest rate on a bank deposit, the rate of return on government bonds, etc.

"NPV=-IC+\Sigma\frac{CFn}{(1+r)^n}=-10000-10000+\frac{6500}{(1+0.15)^1}+\frac{6500}{(1+0.15)^2}\frac{6500}{(1+0.15)^3}=-5159.04"


Internal Rate of Return (IRR, internal rate of return, internal rate of return, internal rate of return, internal rate of return, internal discount rate, internal efficiency coefficient, internal payback coefficient) – a coefficient that shows the maximum allowable risk for an investment project or the minimum acceptable level of profitability. The internal rate of return is equal to the discount rate at which there is no net discounted income, that is, zero

"0=-IC+\Sigma\frac{CFn}{(1+r)^n}"

You can quickly calculate the IRR using the built-in function: internal rate of return